OpenStudy (anonymous):

Find the equation of the line with the given equation" SLOPE -3, through (-2,7)

OpenStudy (anonymous):

u have any idea, how to solve this problem?

OpenStudy (anonymous):

not yet. i'm breaking it down from the book right now.

OpenStudy (anonymous):

okay...u dont know to write a equation of straight line in any form?

OpenStudy (anonymous):

like...y= mx+c

OpenStudy (anonymous):

this is called slope-intercept form.

OpenStudy (anonymous):

u r familiar with this?

OpenStudy (anonymous):

I was just working on that concept (all within the last hour or so) and it is slowly sinking in. But, no, it i snot something I am completely familiar with.. yet... give me a little time with it and I am confident I will conquer it.

OpenStudy (anonymous):

okay...as u have mentioned a word 'slope' in ur question? what does it represent?

OpenStudy (anonymous):

Would it look something like this: -2 = -3x + 7

OpenStudy (anonymous):



OpenStudy (anonymous):

The slope is the rise and run

OpenStudy (anonymous):

what does in mathematically represent?

OpenStudy (anonymous):



OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

Woe... I have never heard of those words. Theta? I don't see that in the book anywhere. I'm working out of the Math94 4th addition book. Chapter 3

OpenStudy (anonymous):

u have heard trigonometrical ratios....sin, cos, tan?

OpenStudy (anonymous):

or if u want answer only. then suppose u have been given slope of the line is 'm' and one point on that line is(x1,y1) then equation of line is: y- y1= m(x-x1)

OpenStudy (anonymous):

No. It has been more then 20 years since I was in a math class. I am in the per-algebra class at the college

OpenStudy (anonymous):

wow...:)

OpenStudy (anonymous):

Would it look something like this: -2 = -3x + 7 I don't want answers only, I want to understand... but I need everything broken down to elementary level because my brain has not ben in 'math mode' in more then 2 decades.

OpenStudy (anonymous):

suppose u have been given slope of the line is 'm' and one point on that line is(x1,y1) then equation of line is: y- y1= m(x-x1) using this u can find the answer of your question.

OpenStudy (anonymous):

any purpose behind enrolling for maths at college?

OpenStudy (anonymous):

after two decades*

OpenStudy (anonymous):

I have a Religious Studies and Theology degree, but I can't do to much with that degree, so i decided to go back to school and pursue a career in education (where I can teach to the degree I already have) but my math is 'out-dated'... so here I am... updating.. lol... and hopefully, in a year, I will be teaching at the high-school. What about you? How did you get involved in deep thinking math? OK, so lets start this over... cause I am lost... the information given is: Slope -3, through (-2,7).

OpenStudy (anonymous):

but u shud understand the meaning of slope..I mean u shud 1st understand d question completely.

OpenStudy (anonymous):

what is slope?

OpenStudy (anonymous):

for now, we forget sin, cos tan..

OpenStudy (anonymous):

Slope is a ratio that compares the vertical change with the horizontal change. (the incline or decline)